Kai must buy canvas to make a sail for his sailboat. The sail will have the shape of a right triangle. It will be as tall as the mast(the vertical pole) and as wide at the bottom as the boom(the horizontal pole). If the mast measure 18 meters and the boom measure 9 meters, about how much canvas does Kai need to buy?
Answer by stanbon(75887)

Kai must buy canvas to make a sail for his sailboat. The sail will have the shape of a right triangle. It will be as tall as the mast(the vertical pole) and as wide at the bottom as the boom(the horizontal pole). If the mast measure 18 meters and the boom measure 9 meters, about how much canvas does Kai need to buy?
-----
Area = (1/2)(base)(height)
Area = (1/2)(9)(18) = 81 sq meters
===================================
cheers,
Stan H.